Grade 11Mechanics

two balls of different masses m and mb are dropped from two different heights viz a and b the ratio of time taken by the two to drop through these distance

harika
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student, 
Given that the mass of "An" is Ma and 
mass of "B" is Mb. 
them two experience an accelaration of "g" as it is tumbling from a tallness 
introductory speed of both is zero as it is dropped from rest. 
From second condition of movement. 
S = 1/2 a t2 [ since u = 0 , it is not taken. ] 
a = 1/2g t12 
b = 1/2g t22 
a/b = t12/t22 
t1/t2 = √a/b 
this the proportion of times taken by Ma and Mb.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)